Please help me with this problem: Solve A = 1/2 h(b1 + b2) for b2
Found 2 solutions by ankor@dixie-net.com, ptaylor:
Answer by ankor@dixie-net.com(22740)

Solve A = 1/2 h(b1 + b2) for b2
:
Get rid of the fraction, mult eq by 2 and you have:
h(b1 + b2) = 2A
:
Divide both sides by h and you have:
b1 + b2 = 2A/h
:
Subtract b1 from both sides and you have
b2 = (2A/h) - b1

Answer by ptaylor(2198)

Solve A = 1/2 h(b1 + b2) for b2
First lets divide both sides by 1/2h or h/2 (h cannot be zero) and we get:
A/(h/2)=b1+b2
We can show that A/(h/2)=2A/h so we have:
2A/h=b1+b2 subtract b1 from both sides and we get
2A/h-b1=b2 or
b2=(2A-b1h)/h
ck
A=h/2(b1+(2A-b1h)/h)) simplifying we get:
A=h/2(b1h+(2A-b1h))/h or
A=(b1h+2A-b1h)/2
A=2A/2=A

Hope this helps. Happy holidays.---ptaylor